OpenStudy (anonymous):

a piece of charcoal is found to contain 30% of the carbon 14 that it originally had. when did the tree die from which the charcoal came? use 5600 years as the half-life of carbon 14. A(t)=Aoe^kt k<0

OpenStudy (anonymous):

0.3 = e^(kt) where k is given by 0.5 = e^(5600k) Solve for t

OpenStudy (anonymous):

i get a funky answer when i try and solve for k....

OpenStudy (anonymous):

What do you get...

OpenStudy (anonymous):

-1.237......x10^-4

OpenStudy (anonymous):

Yup, that's about right

OpenStudy (anonymous):

oh really? i thought it was supposed to be cleaner than that...

OpenStudy (anonymous):

so that i could plug it back in to get t

OpenStudy (anonymous):

yea, my answer is still way off from the back of the book